a surgery has a success rate of 75%. suppose that the surgery is performed on four patients. what is the probability that the surgery is successful on exactly 2 patients?


1
Expert's answer
2021-07-05T15:17:27-0400

Let "X=" the number of patients who respond yes: "X\sim Bin(n, p)."

Given "n=4, p=0.75, q=1-p=0.25."


"P(X=2)=\dbinom{4}{2}(0.75)^2(0.25)^{4-2}=0.2109375"

The probability that the surgery is successful on exactly 2 patients is "0.211."